I need to find two consecutive whole numbers such that 1/4th of the first number is one more than 1/5th of the second number. I know they are 24 & 25 but I need a formula that will work cause mine isn't working.
Answer by mananth(16946)

I need to find two consecutive whole numbers such that 1/4th of the first number is one more than 1/5th of the second number. I know they are 24 & 25 but I need a formula that will work cause mine isn't working.
let one number be x
the other will be x+1
x/4 -1 = x+1 / 5
x+4/ 4 = x+1 / 5
5(x-4) = 4(x+1)
5x-20=4x+4
5x-4x=20+4
x= 24
x+1 =25